A 5.59-kg ball hangs from the top of a vertical pole by a 2.49-m-long string. The ball is struck, causing it to revolve around the pole at a speed of 4.33 m/s in a horizontal circle with the string remaining taut. Calculate the angle, between 0
Answer #1

Tcosθ = mg = 5.59 x 9.81 = 54.84 ( in vertical)

Tsinθ = mv2 /r = 5.59 x 4.332 /2.49sinθ

Tsin2 θ = 42.10 and sin2 θ = 1 - cos2 θ

T(1 - cos2 θ ) = 42.10

T (1 - (54.84/T)2 ) = 42.10

T2 - 42.10T - 3007.43 = 0

T = 79.79 N

79.79cosθ = 54.84

θ = cos-1 (54.84/79.79) = 46.58 degrees